# Hunting the Ghost: Towards Automatic Mining of IoT Hidden Services

### v1

@@ I-B IoT Firmware
According to previous works [19], embedded-Linux based firmware takes up the largest portion of off-the-shelf firmware with a ratio of around 86%. Contents of an embedded-Linux based firmware vary among different vendors. In most cases, there is at least one embedded file system inside and can be extracted by firmware unpacking tools like binwalk [url_binwalk].
@@ I-D IoT interfaces and services
    Others. There are additional interfaces utilizing different kinds of signals. For example, Google home  responds when it hears legal voice commands. Some smart thermometers commits certain behaviors when the temperature has reached a threshold.
@@ IV Request Processing Analysis
    Some embedded web servers provide APIs to associates a URL with a specific handler. For example, GoAhead  provides an API called websUrlHandlerDefine whose first parameter is a string representing the target URL and the second parameter is a processing function object, whenever the target URL is detected, the corresponding handler will be invoked to process the request. We refer to such functions as request-handler binders since server itself will initiative deliver requests to their handlers without querying a tree-like structure.
@@
Function table recognition. Apart from the tree-like structure, function tables are often used to find corresponding handlers. A function table is often composed of multiple compactly-stacked items. Each item has a key(name) and a value(handler address). Once a key matches that of a request, the corresponding handler will be invoked. Table I shows a real function table we extract from our firmware database and Listing LABEL:xx illustrates how function table is usually used inside a function.
@@ V-A Service Categorization
Referring to attributes of services listed in section LABEL:xxx, we summarize the services into four categories:
@@ V-B Service Recognition
API encapsulation. Another issue we consider is the encapsulation of APIs. According to our observations, dynamically linked shared object libraries may introduce functions with the same capabilities but different names from the APIs listed in table II. For example, we found libputil.so of the firmware D-Link DIR-505 v1.09B02 defines a function called _system, which encapsulates the API system by adding a call to vnsprintf in the front. Two functions behave in the almost same way but are invoked interchangeably in other binaries, as shown in the listing LABEL:. To fix this, IoTBolt scans the whole file system(in the firmware) looking for dynamically linked shared libraries used by the binary being analyzed, if there are any function found to be the encapsulation of the APIs in table II and are invoked inside the current binary, IoTBolt add them to the end points list.
@@ VI Requests Recovery by Symbolic Execution
Solution 2. Our high-level idea of solving challenge 2. is to leverage static analysis to guide the symbolic execution. To be specific, we statically generate a set of chopped control flow graphs for each pair of start point and end point. The graphs are used to guide the system in avoiding unnecessary and unrelated paths during the symbolic execution. The detailed algorithm is introduced in section LABEL:xx. Apart from that, we take the strategy of high-level constraints recording. On one hand, this strategy helps the symbolic execution engine bypass functions that have no or little impacts on request recovery, on the other hand, provide security analysts with clearer information about the request contents rather than byte- or bit-level memory constraints.

### v2

@@ I-B IoT Firmware
According to previous works [22], embedded-Linux based firmware takes up the largest portion of off-the-shelf firmware with a ratio of around 86%. Contents of an embedded-Linux based firmware vary among different vendors. In most cases, there is at least one embedded file system inside and can be extracted by firmware unpacking tools like binwalk [14].
@@ I-D IoT interfaces and services
    Others. There are additional interfaces utilizing different kinds of signals. For example, Google home [15] responds when it hears legal voice commands. Some smart thermometers commits certain behaviors when the temperature has reached a threshold.
@@ IV Request Processing Analysis
    Some embedded web servers provide APIs to associates a URL with a specific handler. For example, GoAhead [7] provides an API called websUrlHandlerDefine whose first parameter is a string representing the target URL and the second parameter is a processing function object, whenever the target URL is detected, the corresponding handler will be invoked to process the request. We refer to such functions as request-handler binders since server itself will initiative deliver requests to their handlers without querying a tree-like structure.
@@
Function table recognition. Apart from the tree-like structure, function tables are often used to find corresponding handlers. A function table is often composed of multiple compactly-stacked items. Each item has a key(name) and a value(handler address). Once a key matches that of a request, the corresponding handler will be invoked. Table I shows a real function table we extract from our firmware database and Listing 1 illustrates how function table is usually used inside a function.
@@ V-A Service Categorization
Referring to attributes of services listed in section IV, we summarize the services into four categories:
@@ V-B Service Recognition
API encapsulation. Another issue we consider is the encapsulation of APIs. According to our observations, dynamically linked shared object libraries may introduce functions with the same capabilities but different names from the APIs listed in table II. For example, we found libputil.so of the firmware D-Link DIR-505 v1.09B02 defines a function called _system, which encapsulates the API system by adding a call to vnsprintf in the front. Two functions behave in the almost same way but are invoked interchangeably in other binaries, as shown in the listing 2. To fix this, IoTBolt scans the whole file system(in the firmware) looking for dynamically linked shared libraries used by the binary being analyzed, if there are any function found to be the encapsulation of the APIs in table II and are invoked inside the current binary, IoTBolt add them to the end points list.
@@ VI Requests Recovery by Symbolic Execution
Solution 2. Our high-level idea of solving challenge 2. is to leverage static analysis to guide the symbolic execution. To be specific, we statically generate a set of chopped control flow graphs for each pair of start point and end point. The graphs are used to guide the system in avoiding unnecessary and unrelated paths during the symbolic execution. The detailed algorithm is introduced in section VI. Apart from that, we take the strategy of high-level constraints recording. On one hand, this strategy helps the symbolic execution engine bypass functions that have no or little impacts on request recovery, on the other hand, provide security analysts with clearer information about the request contents rather than byte- or bit-level memory constraints.